40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 19 18:10:10 2022, Paco, Update, BHD, Part V of BHR upgrade - POX11 path, LO path, and ITMX Oplev DJI_0167.JPGDJI_0169.JPGDJI_0168.JPG
    Reply  Wed Apr 20 08:20:22 2022, Paco, Update, BHD, Part V of BHR upgrade - PR2 weirdness 
       Reply  Wed Apr 20 18:50:03 2022, Koji, Update, BHD, Part V of BHR upgrade - PR2 weirdness Screenshot_2022-04-20_18-46-11.png
    Reply  Thu Apr 21 16:49:01 2022, Paco, Update, BHD, Part V of BHR upgrade - BS Oplev + LO1 offset 
       Reply  Thu Apr 21 18:42:47 2022, Tega, Update, BHD, Part V of BHR upgrade - Align LO and AS beams to BHD BS 
          Reply  Thu Apr 21 20:33:31 2022, Koji, Update, BHD, AS1 UR OSEM problem localized in the chamber Screenshot_from_2022-04-21_20-13-13.png
             Reply  Fri Apr 22 12:03:09 2022, Tega, Update, BHD, AS1 UR OSEM problem localized in the chamber 
                Reply  Fri Apr 22 12:15:08 2022, Anchal, Update, BHD, Cable post installation 
                   Reply  Mon Apr 25 16:57:57 2022, Anchal, Update, BHD, Cable post installation 
Message ID: 16788     Entry time: Tue Apr 19 18:10:10 2022     Reply to this: 16789   16797
Author: Paco 
Type: Update 
Category: BHD 
Subject: Part V of BHR upgrade - POX11 path, LO path, and ITMX Oplev 

[Yuta, Paco]

We set up POX11 beam path from ITMX chamber to the ITMX in-air table. To do this, we first identified the POX reflection on the ITMX chamber, and then steered the POXM1 (in the BSC) by hand until we cleared the viewport. We also checked that the POX beam is centered on POXM1.

We then decided to slide the LO1 YAW to clear the LO beam path, which was otherwise clipping on the PR2 SOS. The slider (DAC limited) range of -25000 counts was barely enough to clear the SOS comfortably and avoid hitting the POXM1. The LO beam is now hitting LO2 mirror, so LO alignment can proceed from BSC and ITMY chamber.

Finally, we aligned the input ITMX Oplev beam to ITMXOL2, then ITMX, then to ITMXOL2, and finally into the ITMX in-air optical table. We took some photos of the Oplev beam (see Attachments) to note their position.

By the end of in-vacuum work there was still some flashing in the arm cavities, but fine alignment is required.


After closing the ITMX Chamber, and BSC, we moved on to center the ITMXOL beam. We accomplished this by using two mirrors instead of one as was previously the case. This relaxed the angle of incidence a little, but we had to change the path and the position of the QPD. The QPD sum reads ~ 6600 counts versus the ~ >8000 counts it read right before the vent. One attempt at closing the OL loop, and the ITMX starting oscillating in YAW (PIT was ok), so we realized that maybe we flipped the order in which the OL1 / OL2 mirror were arranged and so the YAW loop needed to flip its sign. Indeed after changing the C1:SUS-ITMX_OL_YAW_GAIN from -6.0 to 6.0 the OL_YAW loop is stable.

Attachment 1: DJI_0167.JPG  3.132 MB  Uploaded Tue Apr 19 19:29:30 2022  | Hide | Hide all
DJI_0167.JPG
Attachment 2: DJI_0169.JPG  2.645 MB  Uploaded Tue Apr 19 19:29:51 2022  | Hide | Hide all
DJI_0169.JPG
Attachment 3: DJI_0168.JPG  2.786 MB  Uploaded Tue Apr 19 19:32:50 2022  | Hide | Hide all
DJI_0168.JPG
ELOG V3.1.3-